1 litre of oxygen and 1 litre of hydrogen are taken in a vessel of 2-litre capacity at NTP.the gases are made to combine by applying electric sparks .assume that water is formed quantitatively.how many grams of warter are formed?what is the other components present in the vessel and in what weight?if the vessel is now heated to 100degree celcieus ,what will be the pressure inside the vesselin mm of Hg?

Grade:upto college level

1 Answers

India VK
10 years ago
Equal volumes of Oxygen and Hydrogen imply equal moles of both of them. Here Hydrogen is the limiting regent. So all of hydrogen is consumed and converted into water. Moles of Water produced = Moles of Hydrogen consumed = 1/22.4 Mass of water produced = 1/22.4 * 18 = The other component in the vessel is gaseous Oxygen. Volume of unreacted Oxygen = 0.5 liters = 0.5/22.4 moles = 0.5 * 32/22.4 gms = If vessel is heated to 100 degress, then all of liquid water is converted into steam. Moles of steam = 0.5 Assume ideal gas conditions, the composition of the vessel is: 0.5 moles steam and 0.5 moles oxygen present in a 2 liter vessel at 373 Kelvins. Using PV= nRT one can find out P inside the vessel.